OpenStudy (anonymous):

A circus performer is shot horizontally out of a cannon at a speed of 40m/s. The cannon is located 50m above a safety net below. 1) Find the time the performer is in the air before hitting the net.

OpenStudy (anonymous):

Horizontal motion has nothing to do with vertical motion. Therefore, you simply need to find the time it takes the performer to fall 50 m due to the force of gravity. h = 0.5 * g * t^2 50 = 0.5 * 9.8 * t^2 t = 3.2 s